OpenStudy (anonymous):

There are 40 students in the 6th grade at Garden Valley Middle School. Eight of them have freckles. What is the ratio of 6th graders with freckles to the total number of 6th graders? 1:3 5:1 1:5 3:1

OpenStudy (anonymous):

The ratio is out of 40 people, 8 have freckles This is 8 out of 40 people Represented by: 8/40 Now you have to simplify = 2/10 = 1/5 Does that make sense?